OpenStudy (anonymous):

What are the coordinates of the equation y=4x+2 ?

OpenStudy (anonymous):

Make a table:  Basically, you plug those numbers as x and then get the y. Then you graph.

OpenStudy (anonymous):

Say you plug in 1 as x. y=4x+1 y=4(1)+1 y=5 The coordinate would be (1, 5) Do this to other points till you got about 3. Then plot and graph!

OpenStudy (anonymous):

Put x = 0, y = 2, Put y = 0, x = -1/2 So one of the coordinates of this is (-1/2, 2)..